Need assistance with this
Consider the following probability distribution.
X= 5 7 8
P(X=x) 0.4 0.1 0.5
1. Find the expected value for the discrete random variable X
2. Find the varance of X.
3. Find the standard deviation of X.
4. Find P(X>7) Where X is greater than or equal to 7. Could not insert the equal under the greater sign.
Thank You

Answer by stanbon(75887)

Consider the following probability distribution.
X= 5 7 8
P(X=x) 0.4 0.1 0.5
1. Find the expected value for the discrete random variable X
E(X) = 0.4*5 + 0.1*7 + 0.5*8 = 6.7
-------------------------------
2. Find the varance of X.
var(X) = 1.4177^2 = 2.0099
3. Find the standard deviation of X.
std(X) = 1.4177
------------------------------------
4. Find P(X>=7)
P(X = 7 or 8) = 0.1+0.5 = 0.6
---------------------------
Cheers,
Stan H.
---------------